A large playground for the new town square in Celebration, Florida, is in the planning stages. Here are the six activities to be completed.

Activity

    Hours Needed

Immediate Predecessor(s)

Planning(A)

        20

        -

Buying Supplies(B)

        60

       Planning(A)

Digging/Grading(C )

       100

        Planning(A)

Sawing lumber(D)

        30

        Buying Supplies(B)

Placing lumber(E )

        20

        Digging/Grading(C ) and Sawing(D)

Assembly/Painting(F )

     10

        Pacing lumber(E )

a) Develop an activity-on-arrow(AOA) network for this project

b) What is the project completion time?
Answer #1

1)

b) Project completion time is the longest activity path = A-C-E-F = 20+100+20+10 = 150hrs.